### v1

# KAG-Thinker: Teaching Large Language Models to Think with Human-like Reasoning Process

Abstract
In this paper, we introduce KAG-Thinker, a novel human-like reasoning framework built upon a parameter-light large language model (LLM). Our approach enhances the logical coherence and contextual consistency of the thinking process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. This framework simulates human cognitive mechanisms for handling complex problems by establishing a structured thinking process. Continuing the Logical Form guided retrieval and reasoning technology route of KAG v0.7, firstly, it decomposes complex questions into independently solvable sub-problems(also referred to as logical forms) through breadth decomposition, each represented in two equivalent forms-natural language and logical function-and further classified as either Knowledge Retrieval or Reasoning Analysis tasks, with dependencies and variables passing explicitly modeled via logical function interfaces. In the solving process, the Retrieval function is used to perform knowledge retrieval tasks, while the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary model to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving model to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn dialogues to align the model with our structured inference paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that KAG-Thinker outperforms state-of-the-art models on seven benchmark datasets, achieving an average improvement of 4.1%. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
[…]
1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG), with representative works including Lewis et al. (2020), Trivedi et al. (2023), and Li et al. (2025). These models often use the LLM’s input to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns reinforcement learning-based retrieval augmentation generation, exemplified by Jin et al. (2025b), Sun et al. (2025), and Wang et al. (2025b). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unusable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
To address these problems and inspired by how humans solve complex problems, we propose an innovative structured human-like cognitive reasoning framework, as shown in Figure 1. We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external information retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-problem using its own knowledge, we rely on external search results. However, search results are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module that extracts the core relevant information. To facilitate a clearer understanding, we will detail the following key innovations.
Logical Form based Reasoning. Prior RAG approaches Lewis et al. (2020); Trivedi et al. (2023); Li et al. (2025), for both questions and sub-problems, are restricted to plain text representations, thereby preventing them from effectively utilizing high-quality knowledge graphs. To tackle this issue, logical form proposes four operation functions tailored for frequently encountered problems (see Appendix 8). Each operation comprises two components: a natural language segment (Step) and a logical function expression segment (Action). For generic retrievers like E5 and BGE-M3, we can directly employ the content within the logical form’s Step. Conversely, for knowledge graph retrievers, the Action portion of the Logical form can be utilized. Crucially, logical form facilitates the propagation of dependencies among interdependent sub-problems. For instance, consider "Step1: Who is the director of Hit Parade Of 1947? Action1: Retrieval(s=s1:film[‘Hit Parade Of 1947’], p=p1:director, o=o1:director)" followed by "Step2: When did #1 die? Action2: Retrieval(s=o1, p=p2:deathtime, o=o2:deathtime)". The variables and enable the seamless transfer of text and expressions from the outcome of the first planning step. In this manner, the logical form enhances the logic, rigor, and stability of LLM thinking.
Breadth Decomposition and Depth Solving. Direct retrieval often proves insufficient for complex multi-hop questions, failing to provide immediate answers. For example, "Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420?". To address such issues, we have devised a framework based on breadth decomposition and depth solving. At a macroscopic level, the decomposition of complex questions into logical forms is generally straightforward and typically does not necessitate recourse to external knowledge. For instance, the aforementioned example can be broken down into five constituent sub-questions: "Who is the director of Hit Parade Of 1947?", "When did #1 die?", "Who is the director of Khiladi 420?", "When did #3 die?", and "Which film was directed by the director who died first according to #2 and #4?". The resolution of each individual sub-problem, conversely, may entail ranging from to (maximum number of searches) retrieval operations, contingent upon its inherent complexity. This in-depth approach to sub-problem solving enables each sub-problem to assimilate adequate external knowledge, thereby reaching a solvable state.
Knowledge Boundary Determination. Most prior approaches Asai et al. (2023); Gutierrez et al. (2024) initiate retrieval for every question or sub-problem, even when the information resides within the LLM’s inherent knowledge. However, LLMs do not necessitate external lookups for such straightforward queries. To mitigate this redundancy, we design a knowledge boundary determination module that eliminates unnecessary external retrieval operations. Prior to knowledge boundary determination, the LLM first generates answers to the sub-problems. To ensure the precision of this knowledge boundary determination, we employ two strategies: (1) prompt-based confidence assessment and (2) likelihood-based confidence assessment. Only when both conditions are satisfied is considered that the solution to the sub-problem does not require external retrieval, and the answer generated by the LLM is directly adopted.
Focusing and Reasoning. When the LLM’s intrinsic knowledge proves inadequate for resolving a given sub-problem, the necessity for external information retrieval arises. However, there is a large quantity of irrelevant or low-quality external information. Following the acquisition of external data, the model undertakes a rigorous evaluation of its credibility and systematically filters out low-fidelity or irrelevant content. Subsequently, the model synthesizes this curated external information with its internal KB, progressively cultivating a comprehensive understanding of the problem through iterative reasoning cycles and knowledge refinement. This iterative process closely emulates human cognitive mechanisms for knowledge construction, acquisition, and information processing, thus critically underpinning the accuracy and reliability of the final reasoning outcomes.
Ultimately, our research yields a human-like cognitive reasoning framework offering substantial advancements, notably: (1) Facilitating more effective utilization of high-quality KBs, (2) Contributing to a more logically coherent, rigorous, and stable approach to complex problem solving, (3) Enabling autonomous decisions regarding the reliance on intrinsic model knowledge versus external information, thereby reducing unnecessary queries, and (4) Exhibiting a greater capacity to manage noise within retrieved data.
2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025b) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity. Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.
3 Approach
We propose an innovative problem solving framework that simulates a human-like reasoning process, as shown in Figure 2. At its core, this framework uses logical form to enable both broad problem decomposition and in-depth problem solving. By integrating a knowledge boundary determination module, we minimize unnecessary external queries, balancing the LLM’s internal knowledge with external information. For sub-problems that require external information, a dedicated focusing and reasoning module is employed to reduce noise in the retrieved content and extract essential information.
3.1 Breadth Decomposition and Depth Solving
Complex multi-hop problems usually need to be broken down to be solved. The key question is how to break them down effectively to best use our existing KBs. We have created logical forms to represent sub-problems. Our method decomposes problems into two parts: broad decomposition, which ensures the main problem and sub-problems stay logical and precise; and in-depth solving, which makes sure Retrieval sub-problem gets enough knowledge to be solved.
3.1.1 Breadth Decomposition
We decompose complex problems breadth-wise into atomic granularity sub-problems, with our decomposition template detailed in Table 1. We define four logical form functions (Retrieval, Math, Deduce, and Output), each dedicated to handling specific tasks: retrieval-focused problems, mathematical computation and causal reasoning tasks, and results aggregation functions. Detailed specifications for these modules can be found in Appendix 8. As shown in Table 1, our breadth decomposition divides the original question into five logical forms of atomic granularity, each independently solvable. The dependency variables are propagated between logical forms by function variable , , and . This approach ensures logically coherent question-solving while maintaining consistent sub-question granularity. To ensure compatibility with both natural language and Logical Form retrievers during sub-question solving, each sub-question adopts dual representations, Step and Action, that maintain semantic equivalence.
| You are an expert in function calls, capable of accurately understanding function definitions and precisely decompose user queries to select appropriate functions to solve problems. The functions are as follows: |
| Function Name: Retrieval |
| Description: Search for SPO information. S stands for subject, O stands for object, represented as variable_name… |
| Function Usage: Retrieval(s=s_alias:type[‘name’], p=p_alias:edge, o=o_alias:type[name], p.prop=‘value’, s.prop=‘value’, o.prop=‘value’) |
| Function Name: Math |
| Description: Perform calculations, which include set operations such as numerical calculations or sorting… |
| Function Usage: Math(content=[‘known conditions’ or ‘o_alias/s_alias’], target=‘goal to be solved’)->math_alias |
| Function Name: Deduce |
| Description: Inference refers to the process of inferring search or calculation results to answer questions… |
| Function Usage: Deduce(op=judgement|entailment|extract|choice|multiChoice, content=[‘known conditions’ or ‘o_alias/s_alias’], target=‘goal to be solved’)->deduce_alias |
| Function Name: Output |
| Description: Directly output A, B, … as the answer, where A and B are variable names referencing previous… |
| Function Usage: Output(A,B,…) |
| Please, based on the definition of the above function, decompose the user question into one or multiple logical steps, outputting the execution plan for each step along with the corresponding action. Please note: Step: Accurately point out the logical thinking process of the question, and use #1 to refer to the solution result of Step1, #2 to refer to the solution result of Step2, and so on Action: Indicate exactly the function you selected and its parameters. |
| Question: Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420? |
| Output: <think>This is a comparison problem. To address the question of which film, between "Hit Parade Of 1947" and "Khiladi 420" was directed by the director who passed away first, we will undertake a chronological investigation. Our initial step involves identifying the director of "Hit Parade Of 1947". Following this, we will ascertain the director of "Khiladi 420". With both directors’ identities established, we will then determine which of these individuals died earlier. This chronological analysis will enable us to connect the director’s lifespan to the films they directed, ultimately revealing which film was associated with the director who died first. By meticulously comparing the timelines of the directors’ lives, we ensure that our conclusion is based on accurate historical data.</think>nn<answer>Step1: Who is the director of Hit Parade Of 1947?nAction1: Retrieval(s=s1:film[‘Hit Parade Of 1947’], p=p1:director, o=o1:director)nStep2: When did #1 die?nAction2: Retrieval(s=o1, p=p2:deathtime, o=o2:deathtime)nStep3: Who is the director of Khiladi 420?nAction3: Retrieval(s=s3:film[‘Khiladi 420’], p=p3:director, o=o3:director)nStep4: When did #3 die?nAction4: Retrieval(s=o3, p=p4:deathtime, o=o4:deathtime)nStep5: Which film was directed by the director who died first according to #2 and #4?nAction5: Deduce(op=choice, content=[‘o2’,‘o4’], target=‘which film was directed by the director who died first according to o2 and o4’)->o5 </answer> |
3.1.2 Depth Solving
Constrained by sub-problem representations and retriever capabilities, many sub-problems defy resolution through single-turn retrievals. We have engineered a depth-oriented resolution framework for Retrieval sub-problem, as detailed in the right part of Figure 2. During depth solving, we prompt the LLM to iteratively search across multi-level and multi-dimensional contexts. The iterative process continues until one of the following conditions is met: (1) the maximum number of turns is reached, or (2) the model generates a conclusive response enclosed between the designated answer tokens <answer> and </answer>. Before conducting depth solving of the Retrieval sub-problem, we first perform a knowledge boundary determination on the sub-problem to assess whether the LLM can directly answer it. If the LLM can confidently answer the current sub-problem, there is no need for depth solving, and the answer can be directly obtained; otherwise, an depth solving process will be initiated. After retrieving relevant content, we conduct focusing and reasoning analysis, allowing the model to determine contextually whether to initiate the next action. If a subsequent action is required, a new Logical Form is generated; otherwise, the sub-question answer is produced. This process iterates interactively until either the maximum number of turns is reached or the sub-question answer is obtained.
3.2 Knowledge Boundary Determination
When addressing complex problems, LLMs commonly decompose them into a series of constituent sub-problems. For the resolution of these sub-problems, two principal strategies are conventionally employed: (1) direct leveraging of the model’s intrinsic KB, and (2) external information retrieval to augment the solution’s correctness and scope. However, each of these approaches presents distinct inherent limitations:
Limitations of Internal Knowledge Utilization. The training paradigm based on maximum likelihood estimation during the pre-training phase induces a strong generative propensity within the model’s parameter space. Consequently, when confronted with ambiguous or unknown information, the model can generate erroneous or fabricated responses. This phenomenon, widely recognized as the hallucination problem, stems from the model’s unwarranted epistemic overconfidence and significantly compromises the veracity of the final output.
Limitations of External Knowledge Retrieval. External knowledge retrieval is indeed capable of mitigating limitations stemming from a model’s intrinsic KB. However, indiscriminate retrieval across all sub-problems engenders significant inefficiencies, manifesting as superfluous computational expenditure and redundant acquisition of information already encoded within the model’s parameters.
Therefore, our work is mainly focused on reasonably determining the knowledge boundary of the model. To achieve this, we propose a Generate First, Then Assess strategy. Specifically, the model first attempts to directly generate answer to the sub-problem. The prompt is shown in Table 2.
| Can you answer the following question step by step? |
[…]
| If you can’t, just reply that based on my internal knowledge, I can’t answer this question, I need to retrieve external knowledge. |
| Question: Who is the author of The Pequod? |
Next, we combine two confidence assessment methods: prompt-based confidence assessment and likelihood-based confidence assessment, to assess the reliability of the generated answer. The advantage of this approach is that it allows the model to first try to use the internal KB to answer the question. For a sub-problem with high enough confidence in the generated answer, this generated answer will be directly used as the final answer. Otherwise, the final answer will be obtained with the help of external knowledge.
1) Prompt-based Confidence Assessment. This confidence assessment method involves gauging confidence by asking the model sub-problems and assessing the authenticity of its generated answer. Specifically, after the model answers a sub-problem, the sub-problem and the generated answer are fed back into the model, prompting it to determine whether they match. Through this prompt-based interactive method, the model can leverage its internal knowledge to reflect and verify its generated answer step by step, ultimately generating a judgment of True or False. The details of the prompt are shown in Table 3.
| Analyze whether the answer to the following question is correct? |
| Please think step-by-step before arriving at your conclusion. |
| If yes, answer True; if no, answer False. |
[…]
| Question: Who is the author of The Pequod? |
| Answer: Herman Melville |
[…]
Once the confidence score is obtained, we can introduce a predefined threshold to further determine whether the generated answer is considered reliable. If the confidence score satisfies: , then the confidence of the generated answer is considered True; otherwise, it is considered False.
| According to the question and references, analyze the relationship between each reference and the question, then summarize whether these references can answer the above questions and give reasons. Wrap the answer and the corresponding reasons in <answer>Yes/No</answer> and <reason>your reason</reason>. |
| References: |
| 0. "Splash (U.S. TV series)" Splash (U.S. TV series) Splash is an American competition-based reality show with celebrity diving competitions which broadcast on ABC from March 19 to May 7, 2013. It was based on the Dutch reality franchise ""Celebrity Splash!"" created by Eyeworks for the series ""Stars Jumping on Saturday"" that premiered in 2012. The show was hosted by actor Joey Lawrence and sportscaster Charissa Thompson with former Olympic divers Steve Foley from Australia and David Boudia from the U.S. as judges. Greg Louganis, also a former U.S. Olympic diver, was the diving instructor and mentor to the celebrities. The series was first |
| 1. "Celebrity Splash!" Celebrity Splash! Celebrity Splash! is a reality television franchise created by Dutch company Eyeworks, started from their Dutch reality show ""Sterren Springen Op Zaterdag"" which premièred in 2012. The franchise involves celebrities diving into the pool. ""Splash"" has its origin and idea from German Olympic-themed variety TV show """" (""TV total Diving""), it was first aired on 16 December 2004, in the ""TV total"" show, on ProSieben and was founded by Stefan Raab and hosted by Sonya Kraus. Other hosts/reporters include Ingolf Lück (2004), Kai Pflaume (2005), Oliver Welke and (2007, 09), Steven Gätjen (2011–12) and (2011–12 reporter). It has |
| 2. "Celebrity Splash! (Argentinian TV series)" Celebrity Splash! (Argentinian TV series) Splash! is a reality television show which teaches celebrities the art of diving. The first series is featured to start on 11 June 2013, and it will be broadcast by Telefe. Marley Wiebe will be hosting the show, with Pampita Ardohain, Maximiliano Guerra, Miguel Ángel Rodríguez and Mariana Montes as judges. ""Splash"" has its origin and idea from German TV Olympic-themed variety show (""TV Total Diving""), it was first aired on 16 December 2004, in the TV total slot, on ProSieben and was founded by Stefan Raab and hosted by Sonya Kraus. Other hosts/reporters |
| Question: Who is the host of Splash! |
| Output: |
| </answer>Yes</answer> |
| <reason>The host of the U.S. TV series Splash is actor Joey Lawrence, alongside sportscaster Charissa Thompson. This is explicitly confirmed in reference 0, which details the U.S. version of the show. References 1 and 2 discuss the Dutch and Argentinian iterations of the franchise, which have different hosts (e.g., Sonya Kraus, Marley Wiebe) and are unrelated to the U.S. version.</reason> |
After obtaining the confidence results from the two assessment methods, we adopt a dual validation strategy to combine them. Specifically, the final confidence is judged as True only when both confidence assessment methods independently yield True. This indicates that the model’s generated answer is reliable and can be directly utilized. Conversely, if either of the two assessment methods results in False, the final confidence is deemed False. In this case, the model’s generated answer is considered potentially unreliable and requires external knowledge to obtain a more accurate answer.
3.3 Focusing and Reasoning
During knowledge boundary determination, we ascertain whether a sub-question requires retrieval. After executing retrieval, we need to assess if the current retrieval results can adequately address the sub-problem. Therefore, we have designed Focusing and Reasoning modules to fulfill this purpose. Focusing and Reasoning are primarily designed to analyze reference materials, with their data construction methodology as detailed in Table 4. Subsequently, based on the answer and reason from the Output section of Table 4, we employ LLMs to perform conditional restructuring, ultimately embedding the restructured content into the designated <think> within the depth-solving module.
3.4 Logical Form Solver
Logical Form, through its dual representation, not only ensures more rigorous logical propagation between sub-problems but also accommodates both natural language retrievers and logical form solvers. The retrieval and reasoning of sub-questions are implemented by Logical Form Solver in KAG framework, which consists of four types of logical form executors: Retrieval Executor is used to retrieve knowledge from an LLM’s internal KB or external KB. Math Executor and Deduce Executor are utilized for Reasoning Analysis. Math is used for math or code calculations, and Deduce is an LLM causal reasoning task. Output Executor is used to determine and output the answer to the question. These four executors are used to connect LLMs with external executors, transforming natural language sub-problems into a structured problem-solving process. The grammatical structure of the logical form is explained in Appendix 8. Table 5 shows an example of using these four executors to solve the problem of calculating the amount of credit card fraud crime. During the inference phase, KAG-Thinker generates equivalent representations of natural language and logical function for each logical form (such as Step1 and Action1 in table 5), and then calls the corresponding Executor based on the dependencies.
| Question: Zhang found a wallet on the street, which contained 5,000 yuan in cash and a credit card. Zhang took the cash and used the credit card to make purchases at four stores: A, B, C, and D, for 210.4 yuan, 569.2 yuan, 1035.2 yuan, 2044.5 yuan, and 1035 yuan, respectiely. After the incident, the public security authorities determined that Zhang was suspected of embezzlement and credit card fraud. What is the total amount involved in the credit card fraud charge against Zhang? | ||
| Steps | Actions | Result |
| 1. What’s the elements of the offense credit card fraud? | Retrieval(s=s1:offense[credit card fraud], p=p1: found, o=o1:elements) | o1 = “The elements of the offense of credit card fraud include the following points: 1….” |
| 2. Which amounts are considered part of the credit card fraud amount? | Deduce(op=extract, content=[o1, Zhang found a wallet on the street…], target=Which amounts are considered part of the credit card fraud amount?)->o2 | o2 = “The total amount involved in Zhang’s credit card fraud is 210. 4 yuan, 569.2 yuan, 1035.3 yuan, 2044.5 yuan, and 1035 yuan.” |
| 3. What’s the total amount involved in Zhang’s credit card fraud? | Math(content=[o1, o2], target=What’s the total amount involved in Zhang’s credit card fraud.)->o3 | o3 = “the total amount involved in Zhang’s credit card fraud is 4989.40 yuan.” |
| 4. Output #3. | Output(o3) | Amount: 4894.40 yuan. |
Retrieval. The logical function of Retrieval is defined as (, , , .contraints, .constraints, .constraints). The function variables are composed in the form of Subject-Predicate-Object (SPO) (, , ) together with conditional constraints. For example, "Who was the director of the 2002 film Men in Black?", which can be transformed into (: Movie[Men in Black], : DirectedBy, : Person, .ReleaseYear=2002). The constraints can be applied to , , or . The final retrieval results include the subgraph and chunks related to target variable object .
Conventional RAG frameworks primarily employ Dense Passage Retrieval (DPR) Karpukhin et al. (2020); Reichman and Heck (2024) based on text similarity by dense vectors. However, semantic similarity is not equivalent to content relevance; retrieving information that is semantically close but irrelevant to the query will harm the reliability of the output. Furthermore, unstructured text corpora introduce different aspects of a topic or introduce multiple events in the same passage, posing a significant challenge for conventional DPR in identifying specific support facts or multi-hop answers where pertinent information may be dispersed or semantically diluted. For example, to solve the problem "Do Bloody Mary and Sidecar share any ingredients in common?", when querying the sub-question "What the ingredients of Bloody Mary?", DPR can easily match different entities named "Bloody Mary" in the corpus, such as cocktail, people, legends, and music. As one of the most complex cocktails, the knowledge of the ingredients of Bloody Mary of regular or specific version is scattered in passages. If there is a purer propositional representation of this knowledge, it can help information focusing when generating answers.
To address these limitations, we continue the KAG’s Text-Graph mutual indexing technology route and upgraded the Hybrid Graph Retrieval method. First, we leverage LLMs to perform Open Information Extraction (OpenIE) and knowledge semantic alignment Liang et al. (2024), construct graph structure index for each document. We define different node types and their relationships to extract and represent knowledge contained in corpus at different semantic granularity. For example, Chunk represents a segment of the original text content; Knowledge Units are independent propositions extracted from the original text. They are used to describe simple fact or event, convey a problem-solving process, algorithm, or principle, as well as cases of a specific domain. As the basis of the KB, knowledge units refine the core information of chunks and are related to each other through Entities and Semantic concepts. In addition, Atomic Query summarizes the questions that can be answered by a knowledge unit to facilitate locating the support facts to answer a sub-question from corpus. For detailed implementation, see Appendix 9.
Second, as shown in Figure 4, we build a hybrid graph retriever in KAG to retrieve this structured index. This hybrid retrieval method consists mainly of five steps: Step 1. DPR: Use the natural language Step part in the specific logical form to retrieve the top-k chunks by matching the name or content with vector similarity. Step 2. Seed Node Focusing: Identify initial relevant knowledge units, atomic queries, and entities of the sub-problem as seed nodes for Personalized PageRank (PPR). The key steps are shown in Figure 5. Step 3. PPR: The top-k entities, top-k knowledge units, and top-k atomic queries retrieved in Step 2 are integrated as seed nodes, and PPR Yang et al. (2024b) runs on the entire graph. Identify the top-k chunks through probability propagation on the path. Step 4. Rerank: For the chunks obtained by DPR and PPR, top-k chunks are selected, respectively, and then rerank them through Reciprocal Rank Fusion (RRF) Cormack et al. (2009). Step 5. Collect Facts: The top-k chunks and the top-k SPOs matched by the logical form are sent to KAG-Thinker as supporting facts for answer generation in this step.
Deduce. The logical function of Deduce is defined as (op, content, target) , where op specifies the reasoning method, including extract, judgement, entailment, and choice. The content field specifies the reasoning context, which includes the information related to the current sub-problem extracted from the user question and the calculation result of the previous logical forms passed through the intermediate variable. The target represents the current objective of the Deduce operation, and represents the result variable of the Deduce reasoning.
Math. The logical function of Math is defined as (content, target) , where content and target are similar to Deduce, specifying the known conditions required for the calculation and the computation task, respectively. represents the result of the computation.
Output. When the question can be answered using either the model’s internal knowledge or the output of previous executors, the output executor will then determine the format and deliver the final answer.
4 Experiments
4.1 Experimental Settings
Benchmarks. Our experiments are conducted on 7 widely-used datasets. The dataset comprises two primary categories: (1) Single-Hop QA: NQ Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), and PopQA Mallen et al. (2023). (2) Multi-Hop QA: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), Musique Trivedi et al. (2022), and Bamboogle Press et al. (2023). These datasets encompass diverse search and reasoning challenges, thereby enabling comprehensive evaluation of our model. Our evaluation set maintains methodological consistency with established prior works Chen et al. (2025); Jin et al. (2025b).
Evaluation Metric. While EM Yu et al. (2024) provides strict verification of answer precision; however,its binary nature is overly restrictive for open-ended scenarios where semantically equivalent answers may exist legal phrasing differences. The F1 metric therefore serves as a nuanced supplement, particularly valuable when evaluating free-form textual output. Comparison Methods. To systematically assess KAGThinker’s effectiveness, we establish a three-tiered evaluation framework comprising the following state-of-the-art baselines: (1) Non-Retrieval Paradigms Naive Generation: Direct answer synthesis without external knowledge integration; Chain-of-Thought (CoT): Explicit cognitive pathway formalization through sequential reasoning traces Wei et al. (2022). (2) Retrieval-Augmented Architectures Naive RAG: Standard retrieval-generation pipeline without iterative optimization Lewis et al. (2020); IRCoT: Multi-cycle retrieval-reasoning coordination with dynamic feedback mechanisms Trivedi et al. (2023); Search-o1: Agent-mediated search workflow integration in reasoning processes Li et al. (2025). (3) Reinforcement Learning Frameworks R1-Gen: Pure RL-optimized generation without search engine interfacing DeepSeek-AI et al. (2025); Search-R1: Multimodal trajectory optimization combining search interactions with reasoning steps Jin et al. (2025b); ZeroSearch: Supervised LLM transformation into dual-function retrieval module (relevant/noisy document generation) Sun et al. (2025); StepSearch: PPO-based search LLM training with incremental exploration Wang et al. (2025b).
Implementation Details. To maintain consistency with previous work, we selected NQ and HotpotQA as our training datasets. Using the data construction and evaluation frameworks outlined in Appendices 10 and 11, we obtained a total of 71K training samples. All comparison methods use Qwen2.5-7B-Instruct as the base model. We employ E5-base-v2 Wang et al. (2024) as the retriever and utilize Wikipedia data from December 2018 as the knowledge base Karpukhin et al. (2020). All corpus indexing and embedding processes are pre-processed using FlashRAG Jin et al. (2025c). Except for ReSearch, which retrieves the top 5 documents for each question, all other methods retrieve the top 3 documents. For additional implementation details, please refer to Appendix 12.
4.2 Main Results
We evaluated our framework on seven widely used benchmark datasets. Here, to maintain consistency with the retriever components of other baseline methods, we conduct plain text retrieval using the "Step" content from the logical form. Table 6 presents the comparisons between our model and other baselines. Compared to baselines without retrieval, the performance of our model exceeds Naive Generation and CoT by an average of 27. 1% and 34. 6%, respectively, in seven datasets. The primary reason for the improvement is that, within these datasets, retrieval is essential, as LLMs struggle to directly answer these questions accurately without access to relevant information. Compared to retrieval-augmented methods, our model outperforms Search-o1, IRCoT, and Naive RAG by average margins of 24.6%, 22.6%, and 14.8%, respectively. These enhancements are primarily attributed to our model’s implementation of breadth decomposition and depth solving, enabling it to learn how to leverage the retriever more effectively—particularly during deep retrieval, where the generated queries align well with the retriever’s capabilities. In comparison to reinforcement learning-based approaches, it can be observed that our model outperforms the previous state-of-the-art model, ReSearch, by an average of 4.1% in EM score across seven datasets. Specifically, it improves by an average of 4.5% on single-hop datasets and by an average of 3.9% on multi-hop datasets. The primary reason is that our breadth decomposition and depth solving framework enables the decomposition of questions into atomic granularity, thereby reducing the complexity involved in model retrieval and answering.
| Single-Hop QA | Multi-Hop QA | Avg | ||||||
| NQ† | TriviaQA∗ | PopQA∗ | HotpotQA† | 2Wiki∗ | MuSiQue∗ | Bamboogle∗ | ||
| Naive Generation | 0.134 | 0.408 | 0.140 | 0.183 | 0.250 | 0.031 | 0.120 | 0.181 |
| CoT | 0.048 | 0.185 | 0.054 | 0.092 | 0.111 | 0.022 | 0.232 | 0.106 |
| Search-o1 | 0.151 | 0.443 | 0.131 | 0.187 | 0.176 | 0.058 | 0.296 | 0.206 |
| IRCoT | 0.224 | 0.478 | 0.301 | 0.133 | 0.149 | 0.072 | 0.224 | 0.226 |
| Naive RAG | 0.349 | 0.585 | 0.392 | 0.299 | 0.235 | 0.058 | 0.208 | 0.304 |
| R1-Gen | 0.270 | 0.537 | 0.199 | 0.237 | 0.292 | 0.072 | 0.293 | 0.271 |
| Search-R1 | 0.393 | 0.610 | 0.397 | 0.370 | 0.414 | 0.146 | 0.368 | 0.385 |
| ZeroSearch | 0.436 | 0.652 | 0.488 | 0.346 | 0.352 | 0.184 | 0.278 | 0.391 |
| [2pt/2pt] StepSearch | - | - | - | 0.386 | 0.366 | 0.226 | 0.400 | - |
| ReSearch | 0.407 | 0.611 | 0.423 | 0.419 | 0.412 | 0.205 | 0.400 | 0.411 |
| Thinker | 0.450 | 0.642 | 0.484 | 0.421 | 0.469 | 0.221 | 0.480 | 0.452 |
4.3 Thinker with KAG
To maintain consistency with KAG-V0.7 Liang et al. (2024), we continue to use its test set, with 1,000 examples each from HotpotQA, Musique, and 2Wiki (consistent with HippoRAG Gutierrez et al. (2024)). The data for our self-built corpus are entirely derived from the supporting facts of these three datasets, totaling 26,990 documents. Within the self-built corpus, and utilizing the same retriever and retrieved documents, our Thinker model still outperforms ReSearch and Search-R1 across three multi-hop datasets. For a fair comparison against the baselines, the original Thinker model could only perform pure text retrieval with BGE-M3 for natural language steps in its logical form, even for operations like Deduce and Math. However, the KAG framework robustly supports these operations. Table 7 illustrates that KAG-Thinker 7B, operating within the KAG framework, shows a marked improvement over Thinker, with average EM and F1 scores increasing by 3.0% and 3.8%, respectively. These enhancements are mainly attributable to KAG-Thinker 7B’s integration of KAG’s hybrid retriever (Sec 3.4) and KAG’s native support for Math and Deduce. This advancement allows KAG-Thinker 7B to achieve performance comparable to the larger KAG-V0.7 72B model, significantly cutting down on online inference demands and resource consumption.
| Framework | Retriever | HotpotQA | MusiQue | 2Wiki | Avg | ||||
| EM | F1 | EM | F1 | EM | F1 | EM | F1 | ||
| Search-R1 | BGE-M3 | 0.545 | 0.676 | 0.307 | 0.403 | 0.517 | 0.589 | 0.456 | 0.556 |
| ReSearch | BGE-M3 | 0.538 | 0.671 | 0.322 | 0.423 | 0.555 | 0.633 | 0.472 | 0.576 |
| Thinker | BGE-M3 | 0.538 | 0.664 | 0.337 | 0.442 | 0.596 | 0.657 | 0.490 | 0.588 |
| KAG-V0.7 72B | KAG Hybrid Retriever | 0.574 | 0.706 | 0.326 | 0.458 | 0.657 | 0.741 | 0.519 | 0.635 |
| KAG-V0.7 32B | KAG Hybrid Retriever | 0.545 | 0.672 | 0.271 | 0.383 | 0.642 | 0.724 | 0.486 | 0.593 |
| KAG-V0.7 7B | KAG Hybrid Retriever | 0.443 | 0.557 | 0.181 | 0.276 | 0.419 | 0.501 | 0.348 | 0.445 |
| KAG-Thinker 7B | KAG Hybrid Retriever | 0.568 | 0.698 | 0.350 | 0.469 | 0.644 | 0.711 | 0.520 | 0.626 |
4.4 Stability Analysis
KAG-V0.7 is prone to inconsistencies in its solution steps when a question is solved twice. To mitigate this issue, we developed KAG-Thinker. We incorporate training data for both breadth decomposition and depth solving into KAG-Thinker, teaching LLMs how to effectively decompose and solve questions. With this data integration, KAG-Thinker’s stability markedly improves. For a more intuitive explanation, we introduce a stability score. For evaluation, we ask a question twice using different random seeds, thereby obtaining two distinct solution processes for each question. Then these two outputs are evaluated by LLM. If LLM determines that the two results are consistent, the question’s solution steps are deemed stable; otherwise, they are considered unstable. As illustrated in Figure 6, KAG-Thinker 7B outperforms KAG-V0.7 7B and KAG-V0.7 72B in terms of stability on HotpotQA, 2Wiki, and Musique. Under commonly used temperature parameters of 0.6 and 0.8, KAG-Thinker 7B demonstrates an average improvement of 17.9% and 7.6% over KAG-V0.7 7B and KAG-V0.7 72B, respectively.
5 Medical Field Application
In recent years, LLMs have been widely applied in the medical field. However, their limited interpretability has hindered real-world deployment, especially for complex, multifaceted problems commonly encountered in medical consultations. For instance, disease diagnosis often involves identifying candidate diseases based on symptoms and then performing differential diagnosis by reviewing the clinical manifestations of each candidate disease—a process that benefits from decomposing the problem into simpler sub-problems. Nevertheless, most existing methods attempt direct retrieval or reasoning Liu et al. (2025); Wang et al. (2025a); Tian et al. (2024); Frisoni et al. (2024); Shi et al. (2024), struggling to effectively address such intricate issues .
On the other hand, for relatively straightforward medical commonsense or symptom-related queries, such as "What is the normal range for heart rate?" or "What symptoms are associated with hypoglycemia?", LLMs can often generate satisfactory answers. For more rigorous questions, such as those involving drug composition or dosage analysis, LLMs are prone to hallucinations, making it necessary to retrieve external knowledge. However, existing approaches Frisoni et al. (2024); Shi et al. (2024) typically do not analyze whether retrieval is truly necessary, leading to cases where retrieval introduces additional noise, particularly when the knowledge base is incomplete.
To address these issues, we design an interpretable framework. By decomposing problems into logically dependent sub-problems, the problem-solving logic becomes more rigorous and interpretable. Through knowledge boundary determination, it reduces the introduction of unnecessary noisy references, thereby improving the accuracy of the model’s answers.
| Question: Boy, 3 years old, weighing 16kg, sustained scald burns on both lower limbs from hot water. Physical examination: BP 85/60 mmHg, restless. Both lower limbs (including buttocks) suffered second-degree burns, with a urine output of 15ml/h. The amount of colloid fluid that should be administered in the first 24 hours is ( ). Options: A: 500ml B: 800ml C: 400ml D: 600ml E: 700ml |
| Breadth Decomposition: <think> The user’s question concerns a 3-year-old boy weighing 16kg… First, we need to calculate the fluid volume based on the burn area and the child’s weight…The proportion of colloid fluid to crystalloid fluid is typically 1:1 or 0.5:1, depending on the severity of the burn. Next, we need to clarify several key aspects: … </think> Step 1: Fluid Requirement Formula Action 1: Retrieval(s=s1:Medical formula[‘Fluid Requirement Formula’], p=p1, o=o1) Step 2: Fluid Replacement Formula for the First 24 Hours Action 2: Retrieval(s=s1, p=p2:time, o=o2:time[‘The first 24 hours’]) Step 3: The Percentage of Total Body Surface Area (TBSA) for Both Lower Limbs (Including the Buttocks) Action 3: Retrieval(s=s2:BodyParts[‘Both Lower Limbs (Including the Buttocks)’], p=p3:AreaRatio, o=o3:AreaRatio) Step 4: The Percentage of Total Body Surface Area (TBSA) for Both Lower Limbs (Including the Buttocks) in Children Action 4: Retrieval(s=s2, p=p4:BelongTo, o=o4:Crowd[‘Children’]) Step 5: The Proportion of Colloid Fluid in the Total Fluid Volume for Second-Degree Burns Action 5: Retrieval(s=s3:TreatmentPlan[‘Burn fluid replacement’], p=p5:Colloidal liquid ratio, o=o5:Colloidal liquid ratio, p.Degree of burn=II∘) Step 6: Based on #1, #2, #3, #4, and #5, calculate the amount of colloid fluid to be administered in the first 24 hours. Action 6: Math(content=[‘o1’, ‘o2’, ‘o3’, ‘o4’, ‘o5’, ‘weight=16kg’], target=‘The amount of colloid fluid to be administered in the first 24 hours’)math6 |
| Depth Solving: Subanswer 1: The formula for calculating fluid replacement in burn patients is often based on the Parkland formula… Subanswer 2: According to burn formulas, the basic principle for fluid resuscitation in pediatric burns is: Total fluid volume for the first 24 hours = Body weight (kg) Burned percentage (%TBSA) 1.5 (pediatric coefficient) + baseline physiological fluid requirement. Subanswer 3: The burn area of both lower limbs (including the buttocks) accounts for approximately 46% of the Total Body Surface Area (TBSA) in adults… Subanswer 4: The burned area for both lower limbs (including buttocks) accounts for approximately 46% of TBSA in adults, but the percentage may differ for children and needs to be further verified. <search>Step: The Percentage of Total Body Surface Area (TBSA) for Both Lower Limbs (Including the Buttocks) in Children. Action: Retrieval(s=s2, p=p4:BelongTo, o=o4:Crowd[‘Children’])</search> <Reference>…</Reference> The area of both lower limbs (including the buttocks) for children is calculated as 46% - (12 - age)% = 37%. Subanswer 5: Parkland Formula: 1ml of fluid per kilogram of body weight per 1% of burned area. Half of the total fluid volume is colloid fluid. However, considering the requirements for the first 24 hours and the differences in pediatric cases, I need to confirm further. <search>Step: The Proportion of Colloid Fluid in the Total Fluid Volume for Second-Degree Burns. Action: Retrieval(s=s3:TreatmentPlan[‘Burn fluid replacement’], p=p5:Colloidal liquid ratio, o=o5:Colloidal liquid ratio, p.Degree of burn=II∘)</search> <Reference>…</Reference> According to the 7th Edition of the Surgery Fluid Resuscitation Protocol, the amount of colloid and electrolyte solution to be provided for pediatric patients is 2.0ml, with a ratio of 1:1 between colloid and electrolyte solution. Subanswer 6: <Python>…</Python> 592ml |
| Answer: 1.Determine Burn Area: The patient is …, the total burn area is calculated to be 37%. 2.Calculate Colloid Volume: According to the fluid replacement protocol … Colloid volume = Body weight (kg) × Total burn area (%) × 1 = 16 kg × 37% × 1 = 592 ml. Since 592 ml is approximately 600 ml, option D is selected. |
5.1 Medical Thinker Model
To address the key points mentioned above, within the framework of KAG-Thinker, we use a breadth decomposition approach to decompose the problem into multiple sub-problems and their corresponding logical form, and we also employ knowledge boundary model to determine whether retrieval is necessary. Table 8 illustrates the reasoning process of KAG-Med-Thinker. First, within the breadth decomposition part, the content between <think> and </think> represents the thought process. Here, it is reasoned that calculating the required fluid volume involves considering both the burn area and weight. Consequently, information regarding the percentage of lower limb (including buttock) surface area in children, as well as the ratio of colloidal fluids to crystalloid fluids, should be retrieved for further decomposition of the sub-problems. Secondly, in the depth solving part, each sub-problem is addressed independently. For sub-problems related to retrieval, the system first performs its own reasoning and then evaluates whether the answer is reliable and complete. If not, it generates a retrieval directive using <search>…</search>. The Chunks retrieved for each sub-problem (with top-k=3) are integrated into the context. Since the retrieved Chunks may be lengthy, we perform knowledge compression and information targeting on the results for each sub-problem. The answers are then refined on the basis of the reference. For sub-problems requiring retrieval (e.g., Step 4 and Step 5), a deep retrieval check is employed to assess whether the Reference can resolve the sub-problem. If the Reference cannot address the sub-problem, a new retrieval directive is generated. Specifically, for Step 6, the Math directive is executed by generating Python code and obtaining computational results using a code interpreter. Finally, based on the breadth decomposition and depth solving, the complete reasoning process and the final result are generated.
We structure the samples into a multi-turn dialogue format. For synthesizing data, question decomposition is performed using Qwen2.5-72B-Instruct, which offers stronger reasoning capabilities and instruction-following abilities, allowing distillation into smaller models. The target model is trained to generate answers for sub-problems and determine whether external executors are required. Since the medical Q&A often involves lengthy answers for sub-problems, we call the LLM twice to confirm whether the answers are correct and complete, which serve as labels for the knowledge boundary determination. The subsequent deep retrieval, reasoning, and answer synthesis utilize distilled synthetic data derived from the LLMs.
5.2 Medical Knowledge Graph
5.2.1 Medical Data Sources
We construct a Chinese medical knowledge retrieval repository containing data sourced from authoritative medical textbooks and professional medical websites. The Medical Textbooks originate from the Chinese materials in MedQA, specifically related to resources from the National Medical Licensing Examinations in China (e.g., the Licensed Physician Qualification Examination). These materials are highly professional and practical, making them well-suited for research and development of medical intelligent question-answering systems tailored for Chinese users. Professional Medical Websites. Medical data were primarily collected from five professional medical websites: Xiahe Medical Encyclopedia, DXY Doctor, Tencent Medical Encyclopedia, Kuaishou Medical Encyclopedia, and Baidu Health.
5.2.2 Medical Knowledge Graph Construction
For the extraction of information from the medical domain, we adopt schema constraint extraction due to the low error tolerance in the medical context. The use of a predefined schema restricts the extraction and generation to specific, structured key information (e.g., disease names, drug dosages, treatment plans), avoiding the introduction of irrelevant or erroneous content. We select 23 commonly used medical entity types that cover the main semantic categories involved in clinical diagnosis and treatment (e.g., diseases, symptoms, medications, diagnostic tests) and extend to general concepts related to the medical domain, public health, and system integration (e.g., disciplines, organizations, biological processes). This classification system refers to international medical terminology standards (for example, SNOMED CT, ICD-10) and ontologies of representative medical knowledge graphs, which effectively support subsequent structured knowledge fusion, reasoning, and generation tasks. Specific medical entity types can be found in Appendix 14. The detailed structured knowledge extraction process and prompt references can be found in Appendix 15. Figure 7 illustrates the constructed medical knowledge graph and shows the connectivity of the structured example information.
5.3 Experimental Settings and Results
5.3.1 Experimental Settings
Dataset. We selected MedQA Jin et al. (2020) as the evaluation dataset. This dataset is a collection of question-and-answer pairs sourced from professional medical examinations. All questions are multiple-choice, covering areas such as medical knowledge, disease diagnosis, pharmacology, and diagnostic procedures. Since the original test dataset is very large, we randomly sampled 300 entries from the Chinese subset to form our evaluation dataset.
Implementation Details. We conducted SFT (Supervised Fine-Tuning) training on two base models, Qwen2.5-14B-Instruct and DeepSeek-R1-Distill-Qwen-14B, with the following configurations: learning rate lr=5e-6, epoch=5, and using a self-constructed medical KB in Section 5.2.2, while leveraging BGE-M3 to build vector indexes.
5.3.2 Main Results
The empirical evaluation results are presented in Table 9. KAG-Med-Thinker consistently demonstrates enhanced performance across both base models, with a particularly significant improvement being observed on the DeepSeek-R1-Distill-Qwen-14B. In particular, compared to the established multi-turn planning and retrieval augmented models, IRCoT and ReAct, KAG-Med-Thinker achieves remarkable performance improvements of 3.95%, 4.41% respectively. It also outperforms the adaptive retrieval model, Naive RAG, by a margin of 3.8%. Compared to Qwen2.5-14B-Instruct, KAG-Med-Thinker shows similar superiority, outperforming IRCoT, ReAct, and Naive RAG by 4.37%, 0.8%, and 2%, respectively. These consistent gains across different base models and comparative baselines unequivocally validate the effectiveness and robustness of KAG-Med-Thinker.
| LLM | Framework | MedQA |
| Qwen2.5-14B-Instruct | Naive Generation Qwen et al. (2025) | 84.00 |
| Naive RAG Lewis et al. (2020) | 81.67 | |
| Naive RAG+adaptive Jeong et al. (2024) | 85.00 | |
| IRCoT Trivedi et al. (2023) | 82.33 | |
| ReAct Yao et al. (2023) | 86.20 | |
| KAG-Med-Thinker | 87.00 | |
| DeepSeek-R1-Distill-Qwen-14B | Naive Generation DeepSeek-AI et al. (2025) | 79.67 |
| Naive RAG Lewis et al. (2020) | 79.00 | |
| Naive RAG+adaptive Jeong et al. (2024) | 81.48 | |
| IRCoT Trivedi et al. (2023) | 81.33 | |
| ReAct Yao et al. (2023) | 80.87 | |
| KAG-Med-Thinker | 85.28 |
Furthermore, it is observed that the adaptive retrieval-based Naive RAG consistently outperforms its exhaustive retrieval-based counterpart. In particular, it demonstrates a significant performance advantage of 3.33% on Qwen2.5-14B-Instruct and 2.48% on DeepSeek-R1-Distill-Qwen-14B. This superior performance even extends to IRCoT and ReAct, models that employ multi-turn retrieval planning. This finding shows that unconstrained or unfiltered retrieval introduces inherent limitations and noise that negatively impact overall reasoning performance. Consequently, it is essential to judiciously determine the model’s intrinsic knowledge boundaries and synergistically integrate its internal knowledge with external tool invocation to fully harness their complementary strengths.
6 Conclusion
KAG-Thinker is presented as a novel, cognitively-inspired reasoning framework engineered to address the limitations of extant methodologies in solving intricate multi-hop problems demanding rigorous logical coherence and contextual consistency. The framework’s core methodology emulates human cognitive processes by decomposing complex queries into logically interconnected sub-problems through a structured logical-form-based reasoning mechanism, thereby enabling incremental problem resolution while preserving structural integrity. Each logical-form task is associated with a unique logical function representation and can be solved independently. This design enables the effective decoupling of knowledge retrieval and reasoning analysis into separate subtasks. Consequently, the retrieval component can be seamlessly integrated with an automated knowledge boundary detection module, which optimizes efficiency by intelligently arbitrating between internal knowledge utilization and external information retrieval, thereby enabling the LLM to decide whether to retrieve external information based on its internal assessment. Furthermore, an integrated focusing-and-reasoning module refines retrieved content and synergizes external data with internal inferential processes, mirroring human iterative knowledge refinement. The applicability of KAG-Thinker in high-stakes domains is empirically validated via a medical Q&A system, showcasing its capacity to integrate domain-specific knowledge while maintaining logical rigor.
7 Contributors
Authors are listed alphabetically by the first name.
Dalong Zhang
Jun Xu
Jun Zhou
Lei Liang
Lin Yuan
Ling Zhong
Mengshu Sun
Peilong Zhao
QiWei Wang
Xiaorui Wang
Xinkai Du
YangYang Hou
Yu Ao
ZhaoYang Wang
Zhengke Gui
ZhiYing Yi
Zhongpu Bo