### v2

# KAG-Thinker: Interactive Thinking and Deep Reasoning in LLMs via Knowledge-Augmented Generation

Abstract
In this paper, we introduce KAG-Thinker, which upgrade KAG to a multi-turn interactive thinking and deep reasoning framework powered by a dedicated parameter-light large language model (LLM). Our approach constructs a structured thinking process for solving complex problems, enhancing the the logical coherence and contextual consistency of the reasoning process in question-answering (Q&A) tasks on domain-specific knowledge bases (KBs) within LLMs. Following the Logical Form guided retrieval and reasoning technology route of KAG, this framework first decomposes complex questions into independently solvable sub-problems (which are also referred to as logical forms) through breadth decomposition. Each such logical form is represented in two equivalent forms—natural language and logical function—and subsequently classified as either a Knowledge Retrieval or Reasoning Analysis task. Dependencies and parameter passing between these tasks are explicitly modeled via logical function interfaces. In the solving process, the Retrieval function performs retrieval tasks. It retrieves one-hop structured and unstructured information of specified knowledge unit. While the Math and Deduce functions are used to perform reasoning analysis tasks. Secondly, it is worth noting that, in the Knowledge Retrieval sub-problem tasks, LLMs and external knowledge sources are regarded as equivalent KBs. We use the knowledge boundary module to determine the optimal source using self-regulatory mechanisms such as confidence calibration and reflective reasoning, and use the depth solving module to enhance the comprehensiveness of knowledge acquisition. Finally, instead of utilizing reinforcement learning, we employ supervised fine-tuning with multi-turn interactive thinking and deep reasoning to align the model with our structured reasoning paradigm, thereby avoiding excessive reflection. This is supported by a data evaluation framework and iterative corpus synthesis, which facilitate the generation of detailed reasoning trajectories. Experimental results show that our model outperforms state-of-the-art models on seven benchmark datasets, achieving an average improvement of 4.1%. In addition, KAG-Thinker not only enhances the reasoning stability of the KAG framework but also achieves performance comparable to the model with 10 times its parameters. We further demonstrate its practical effectiveness through a medical Q&A system integrated with domain-specific knowledge. We train KAG-Med-Thinker model with 14b-parameter based on synthetic medical corpora using the same approach, respectively, and verify their applicability in specialized applications.
[…]
1 Introduction
Recent advances in large language models (LLMs) OpenAI et al. (2024); Qwen et al. (2025) have shown impressive performance in Q&A tasks. However, their ability to solve multi-hop problems Wei et al. (2022), especially those that require external retrieval of up-to-date information Jin et al. (2025a), has certain limitations. Previous work primarily focused on how to use external retrievers, broadly falling into two categories: (1) decoupling retrieval from training, and (2) autonomously deciding whether to use the retriever during training. The first category primarily involves retrieval augmentation generation (RAG) with predefined workflow, the representative works include Lewis et al. (2020)Trivedi et al. (2023), and Li et al. (2025a). These models often use the question to retrieve relevant passages, which are then added to the LLM’s context. This helps the LLM leverage external knowledge when answering questions. The second category mainly concerns to mimic the deeper and more deliberative reasoning as "slow thinking"  Pan et al. (2025), a paradigm inspired by human cognitive processes, to determine the timing of retrieval dynamically. Exemplified by Jin et al. (2025b)Sun et al. (2025), and Wang et al. (2025b). These methods generate a specific token <search> that indicates an external search engine query is required. Once the retriever provides results, these are then fed back into the model as input, allowing it to resume content generation. Although these methods improved LLM utilization of retrievers, they do not truly provide a step-by-step approach for solving complex multi-hop problems. Their approaches are more concerned with how LLMs arrive at the correct answer, without ensuring the rigor, stability, or logical consistency of the underlying problem-solving process. Consequently, these methods are often unreliable in high-stakes or sensitive domains such as finance, medicine, and legal applications.
In order to achieve logical reasoning with more accurate judgments and reduced biases (Li et al., 2025c), recent work has attempted various approaches, such as the atomic reasoning actions decomposition by  Wu et al. (2024) , hierarchical and structured reasoning by  Yang et al. (2025) , and verify reasoning trajectories with symbolic solver by  Guan et al. (2025) . Moreover,  Sumers et al. (2024) propose a cognitive architectures for language agents with three key dimensions: information storage, action space and decision-making procedure. Benefit from these previous works and inspired by Slow Thinking - which solves complex problem with access to additional computational resources, full attention, and sophisticated logical reasoning  Booch et al. (2021) - we propose an innovative structured thinking and reasoning model as shown in Figure 1. It enables deep thinking in LLM by simulating human cognitive mechanisms without reinforcement learning DeepSeek-AI et al. (2025); OpenAI (2024). We use logical form to decompose complex problems broadly and solve them in depth. Logical form improves the logic, rigor, and stability of the LLM reasoning process. When humans solve complex problems, they only perform external information retrieval on knowledge that they do not know. Similarly to this process, we do not need to perform external knowledge retrieval for all decomposed sub-problems. Therefore, we design an LLM’s knowledge boundary determination module to reduce unnecessary external retrieval and mitigate noise from the retrieved content. If the LLM cannot confidently answer a certain sub-question using its own knowledge, we rely on external KBs. However, search results from external KB are often unreliable and noisy. To address this, we created a Focusing-and-Reasoning module to extract the core relevant information. To facilitate a clearer understanding, we will introduce these key innovations in detail in the following paragraphs.
Figure 1: Overview of the KAG-Thinker reasoning model. The corpus synthesis pipeline in the figure from left to right: 1) Decompose the original $question$ into multiple sub-questions with clear logical dependencies and can be solved independently, each sub-question corresponds to a unique logical form, which consists of two components. For example, $logicalform_{i}$ consists of $subq\text{\_}NL_{i}$ expressed in natural language and $subq\text{\_}EXP_{i}$ expressed in logical function equivalent. The logical forms are classified into one of $Retrieval$, $Math$, $Deduce$ and $Output$ through function name. 2) Perform knowledge boundary determination and depth solving on $Retrieval$ sub-questions. During depth solving, the focusing and reasoning module reduces noise in retrieved information. 3) Finally, synthesize multi-turn interactive reasoning corpus expressed in natural language and logical function express, and obtain KAG-Thinker Model through supervised fine-tuning of Ling and Qwen’s base and instruct models.
Logical Form based Reasoning. Prior RAG approaches Lewis et al. (2020); Trivedi et al. (2023); Li et al. (2025a), for both questions and sub-questions, are restricted to plain text representations, thereby preventing them from effectively utilizing high-quality structured knowledge. To address this issue, logical form proposes four operation functions tailored to frequently encountered questions (see the Appendix 8). Each operation comprises two components: a natural language segment (Step) and a logical function expression segment (Action). For generic retrievers like E5 and BGE-M3, we can directly employ the content within the logical form’s Step. Conversely, for structured knowledge retrievers, the Action portion of the logical form can be utilized. Crucially, logical form facilitates the propagation of dependencies among interdependent sub-problems. For example, consider "Step1: Who is the director of Hit Parade Of 1947? Action1: Retrieval(s=s1:film[‘Hit Parade Of 1947’], p=p1:director, o=o1:director)" followed by "Step2: When did #1 die? Action2: Retrieval(s=o1, p=p2:deathtime, o=o2:deathtime)". The variables #1\#1# 1 and $o1$ enable seamless transfer of text and expressions from the outcome of the first planning step. In this manner, the logical form enhances the logic, rigor, and stability of the LLM reasoning.
Breadth Decomposition and Depth Solving. Direct retrieval often proves insufficient for complex multi-hop questions, failing to provide immediate answers. For example, "Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420?". To address such issues, we have devised a model based on breadth decomposition and depth solving. At a macroscopic level, the decomposition of complex questions into logical forms is generally straightforward and typically does not necessitate recourse to external knowledge. For instance, the aforementioned example can be broken down into five constituent sub-questions: "Who is the director of Hit Parade Of 1947?", "When did #1 die?", "Who is the director of Khiladi 420?", "When did #3 die?", and "Which film was directed by the director who died first according to #2 and #4?". The resolution of each individual sub-problem, conversely, may entail ranging from 0 to $M$ (maximum number of searches) retrieval operations, contingent upon its inherent complexity. This in-depth approach to sub-problem solving enables each sub-problem to assimilate adequate external knowledge, thereby reaching a solvable state.
Knowledge Boundary Determination. Most prior approaches Asai et al. (2023); Gutierrez et al. (2024) initiate retrieval for every question or sub-question, even when the information resides within the LLM’s inherent knowledge. However, LLMs do not necessitate external lookups for such straightforward queries. To mitigate this redundancy, we design a knowledge boundary determination module that eliminates unnecessary external retrieval operations. Prior to knowledge boundary determination, the LLM first generates answers to the sub-problems. To ensure the precision of this knowledge boundary determination, we employ two strategies: (1) prompt-based confidence assessment and (2) likelihood-based confidence assessment. Only when both conditions are satisfied is considered that the solution to the sub-problem does not require external retrieval, and the answer generated by the LLM is directly adopted.
Focusing and Reasoning. When the LLM’s intrinsic knowledge proves inadequate for resolving a given sub-problem, the necessity for external information retrieval arises. However, there is a large quantity of irrelevant or low-quality external information. Following the acquisition of external data, the model undertakes a rigorous evaluation of its credibility and systematically filters out low-fidelity or irrelevant content. Subsequently, the model synthesizes this curated external information with its internal KB, progressively cultivating a comprehensive understanding of the problem through iterative reasoning cycles and knowledge refinement. This iterative process closely emulates human cognitive mechanisms for knowledge construction, acquisition, and information processing, thus critically underpinning the accuracy and reliability of the final reasoning outcomes.
Ultimately, our research yields a human-like structured iterative thinking and depth solving model offering substantial advancements, notably: (1) Facilitating more effective utilization of high-quality KBs, (2) Contributing to a more logically coherent, rigorous, and stable approach to complex problem solving, (3) Enabling autonomous decisions regarding the reliance on intrinsic model knowledge versus external information, thereby reducing unnecessary queries, and (4) Exhibiting a greater capacity to manage noise within retrieved data.
2 Related Work
Retrieval-Augmented Generation (RAG) integrates relevant, real-time information from external KBs through a retrieval mechanism, thereby enhancing its responses with contextual richness and factual grounding Borgeaud et al. (2022); Dong et al. (2023); Luo et al. (2024). Despite these advantages, a persistent challenge in RAG systems is that retrieved content can often be mixed with irrelevant or erroneous information due to the inherent limitations of current retrieval solutions. Moreover, predefined reasoning adopts structured and modular RAG pipelines that limit flexibility in evolving and open-ended tasks  Li et al. (2025c, b).
To mitigate these challenges, researchers have introduced various components designed to enhance the robustness and effectiveness of RAG systems Gao et al. (2023); Li et al. (2024); Tan et al. (2024). A key focus in recent studies has been on enabling RAG systems to autonomously determine when to leverage retrieval mechanisms. For example, Jin et al. (2025b); Sun et al. (2025); Wang et al. (2025b) introduced a special token to explicitly trigger an external search, with the retrieval results subsequently integrated into the model to support content generation. However, Ren et al. (2025) identified limitations in LLM regarding their ability to accurately distinguish between internal and external knowledge, as well as their difficulty in reconciling conflicts between the two. To address this, they proposed Prior judgment and Poster judgments to evaluate confidence levels and assess response correctness. In parallel, another significant line of research has focused on developing adaptive strategies to optimize how retrieval is utilized, based on query complexity.  Jeong et al. (2024) developed a classifier to intelligently decide between iterative, single, or no retrieval strategies depending on the complexity of the query. Building upon this idea, Islam et al. (2024) proposed a hybrid adaptive retrieval method: The model initially generates an answer without retrieval and then dynamically decides the need for retrieval based on its confidence, with the aim of balancing performance gains with inference speed. Furthermore, Tang et al. (2025) extended this work by applying a reinforcement learning-based framework to select the most appropriate retrieval strategy for a given query. While these diverse approaches have significantly advanced the integration and adaptive use of retrieval mechanisms within RAG systems, they currently lack a comprehensive, step-by-step framework for effectively addressing complex multi-hop reasoning problems. This critical area remains an open challenge for further research.
3 Approach
Figure 2: Overview of the problem-solving model that uses human-like reasoning process. During problem breadth decomposition, all sub-problems are obtained in a single decomposition pass, where each sub-problem is an atomic problem that can be solved independently, also known as a logical form. Herein, the terms Step and Action maintain semantic consistency, both denoting such a sub-problem. Within problem breadth decomposition, Step employs #$\#n$ for answer propagation of the $n$-th sub-problem, while Action binds variables in logical function (e.g., $o_{n}$, $s_{n}$) for variable transmission. By determining the knowledge boundary of the Retrieval sub-problem, it is decided whether to utilize the base model’s answer or to generate a deep retrieval. During depth solving of sub-problems, the system sequentially executes retrieval, focusing, and reasoning in iterative processes until either the sub-problem answer is obtained or the maximum solving attempt threshold is activated.
We propose an innovative complex problem solving model, as shown in Figure 2. At its core, this model uses logical form to enable both broad problem decomposition and in-depth problem solving. By integrating a knowledge boundary determination module, we minimize unnecessary external queries, balancing the LLM’s internal knowledge with external information. For sub-problems that require external information, a dedicated focusing and reasoning module is employed to reduce noise in the retrieved content and extract essential information.
3.1 Breadth Decomposition and Depth Solving
Complex multi-hop problems usually need to be broken down to be solved. The key question is how to break them down effectively to best use our existing KBs. We have created logical forms to represent sub-problems. Our method decomposes problems into two parts: broad decomposition, which ensures the main problem and sub-problems stay logical and precise; and in-depth solving, which makes sure Retrieval sub-problem gets enough knowledge to be solved.
3.1.1 Breadth Decomposition
We decompose complex problems breadth-wise into $n$ atomic granularity sub-problems, with our decomposition template detailed in Table 1. We define four logical form functions (Retrieval, Math, Deduce, and Output), each dedicated to handling specific tasks: retrieval-focused problems, mathematical computation and causal reasoning tasks, and results aggregation functions. Detailed specifications for these modules can be found in Appendix 8. As shown in Table 1, our breadth decomposition divides the original question into five logical forms of atomic granularity, each independently solvable. The dependency variables are propagated between logical forms by function variable #$\#n$, $o_{n}$, and $s_{n}$. This approach ensures logically coherent question-solving while maintaining consistent sub-question granularity. To ensure compatibility with both natural language and Logical Form retrievers during sub-question solving, each sub-question adopts dual representations, Step and Action, that maintain semantic equivalence.
You are an expert in function calls, capable of accurately understanding function definitions and precisely decompose user queries to select appropriate functions to solve problems. The functions are as follows:
Function Name: Retrieval
 Description: Search for SPO information. S stands for subject, O stands for object, represented as variable_name…
 Function Usage: Retrieval(s=s_alias:type[‘name’], p=p_alias:edge, o=o_alias:type[name], p.prop=‘value’, s.prop=‘value’, o.prop=‘value’)
Function Name: Math
 Description: Perform calculations, which include set operations such as numerical calculations or sorting…
 Function Usage: Math(content=[‘known conditions’ or ‘o_alias/s_alias’], target=‘goal to be solved’)->math_alias
Function Name: Deduce
 Description: Inference refers to the process of inferring search or calculation results to answer questions…
 Function Usage: Deduce(op=judgement|entailment|extract|choice|multiChoice, content=[‘known conditions’ or ‘o_alias/s_alias’], target=‘goal to be solved’)->deduce_alias
Function Name: Output
 Description: Directly output A, B, … as the answer, where A and B are variable names referencing previous…
 Function Usage: Output(A,B,…)
Please, based on the definition of the above function, decompose the user question into one or multiple logical steps, outputting the execution plan for each step along with the corresponding action. Please note: Step: Accurately point out the logical thinking process of the question, and use #1 to refer to the solution result of Step1, #2 to refer to the solution result of Step2, and so on Action: Indicate exactly the function you selected and its parameters.
Question: Which film has the director who died first, Hit Parade Of 1947 or Khiladi 420?
[…]
Table 1: Instruction template for breadth decomposition. The template initiates with problem positioning and analysis to determine decomposition steps, ultimately generating all atomic sub-problems through single-pass decomposition. Sub-question dependencies are propagated via #$\#n$ in natural language Steps and referenced through logical function variables (e.g., $o_{n}$, $s_{n}$) in Actions.
3.1.2 Depth Solving
Constrained by sub-problem representations and retriever capabilities, many sub-problems defy resolution through single-turn retrievals. We have engineered a depth-oriented resolution strategy for Retrieval sub-problem, as detailed in the right part of Figure 2. During depth solving, we prompt the LLM to iteratively search across multi-level and multi-dimensional contexts. The iterative process continues until one of the following conditions is met: (1) the maximum number of turns is reached, or (2) the model generates a conclusive response enclosed between the designated answer tokens <answer> and </answer>. Before conducting depth solving of the Retrieval sub-problem, we first perform a knowledge boundary determination on the sub-problem to assess whether the LLM can directly answer it. If the LLM can confidently answer the current sub-problem, there is no need for depth solving, and the answer can be directly obtained; otherwise, an depth solving process will be initiated. After retrieving relevant content, we conduct focusing and reasoning analysis, allowing the model to determine contextually whether to initiate the next action. If a subsequent action is required, a new Logical Form is generated; otherwise, the sub-question answer is produced. This process iterates interactively until either the maximum number of turns is reached or the sub-question answer is obtained.
3.2 Knowledge Boundary Determination
When addressing complex problems, LLMs commonly decompose them into a series of constituent sub-problems. For the resolution of these sub-problems, two principal strategies are conventionally employed: (1) direct leveraging of the model’s intrinsic KB, and (2) external information retrieval to augment the solution’s correctness and scope. However, each of these approaches presents distinct inherent limitations:
Limitations of Internal Knowledge Utilization. The training paradigm based on maximum likelihood estimation during the pre-training phase induces a strong generative propensity within the model’s parameter space. Consequently, when confronted with ambiguous or unknown information, the model can generate erroneous or fabricated responses. This phenomenon, widely recognized as the hallucination problem, stems from the model’s unwarranted epistemic overconfidence and significantly compromises the veracity of the final output.
Limitations of External Knowledge Retrieval. External knowledge retrieval is indeed capable of mitigating limitations stemming from a model’s intrinsic KB. However, indiscriminate retrieval across all sub-problems engenders significant inefficiencies, manifesting as superfluous computational expenditure and redundant acquisition of information already encoded within the model’s parameters.
Therefore, our work is mainly focused on reasonably determining the knowledge boundary of the model. To achieve this, we propose a Generate First, Then Assess strategy. Specifically, the model first attempts to directly generate answer to the sub-problem. The prompt is shown in Table 2.
Can you answer the following question step by step?
[…]
If you can’t, just reply that based on my internal knowledge, I can’t answer this question, I need to retrieve external knowledge.
Question: Who is the author of The Pequod?
Table 2: Instruction template for solving sub-problem.
Next, we combine two confidence assessment methods: prompt-based confidence assessment and likelihood-based confidence assessment, to assess the reliability of the generated answer. The advantage of this approach is that it allows the model to first try to use the internal KB to answer the question. For a sub-problem with high enough confidence in the generated answer, this generated answer will be directly used as the final answer. Otherwise, the final answer will be obtained with the help of external knowledge.
1) Prompt-based Confidence Assessment. This confidence assessment method involves gauging confidence by asking the model sub-problems and assessing the authenticity of its generated answer. Specifically, after the model answers a sub-problem, the sub-problem and the generated answer are fed back into the model, prompting it to determine whether they match. Through this prompt-based interactive method, the model can leverage its internal knowledge to reflect and verify its generated answer step by step, ultimately generating a judgment of True or False. The details of the prompt are shown in Table 3.
Analyze whether the answer to the following question is correct?
Please think step-by-step before arriving at your conclusion.
If yes, answer True; if no, answer False.
[…]
Question: Who is the author of The Pequod?
Answer: Herman Melville
Table 3: Instruction template for confidence assessment. Where Question represents the sub-problem to be solved, and Answer represents the LLM-generated answer.
Figure 3: The probability distribution of the answer tokens. The left shows the low confidence answers, and the right shows the high confidence answers.
According to the question and references, analyze the relationship between each reference and the question, then summarize whether these references can answer the above questions and give reasons. Wrap the answer and the corresponding reasons in <answer>Yes/No</answer> and <reason>your reason</reason>.